# ASURA-FDPS-ML: Star-by-star Galaxy Simulations Accelerated by Surrogate Modeling for Supernova Feedback

### v1

@@ 2.6 Hybrid on-the-fly ML-driven simulations
Figure 1 shows the schematic diagram of our presenting framework, ASURA-FDPS-ML. It has two kinds of MPI communicators: “Main” and “Pool” hereafter. A whole galaxy, including $N$-body and SPH, is directly computed in main processes using hundreds of MPI processes (“Main”), while SN feedback in dense regions ($>1~{}\mathrm{cm^{-3}}$) is handled by a surrogate model in “Pool” processes. As SNe in the less dense ambient gas are more kinetically resolved (Kim & Ostriker, 2015; Steinwandel et al., 2020; Gutcke et al., 2021), when SNe explode in the environments at a hydrogen number density of $<1\mathrm{cm^{-3}}$, the energy of SNe ($10^{51}$ erg) is injected into the neighboring 100 SPH particles weighted using the Wendland $C^{4}$ kernel (Dehnen & Aly, 2012). Every timestep, “Main” checks if any SN occurs in dense regions ($>1~{}\mathrm{cm^{-3}}$) or not. If SNe occur, the ambient gas particles are sent to a “Pool” process. The surrogate model in the “Pool” process reconstructs the distribution of gas particles after a time window $\tau$ from the SN explosion. Once the “Main” processes are advanced for $\tau$, the predicted gas particles are returned to “Main” from “Pool”.
@@ 3 Test for our surrogate model of SN feedback
Figure 2: Comparison between simulations and ML for isolated SN. Upper: The numerical simulation results. Lower: The numerical simulation results with our surrogate SN feedback model. Cross section of snapshots at $t=0,~{}10^{5},~{}\mathrm{and}~{}4\times 10^{5}$ years are listed up from left to right.
Our surrogate model is designed to incorporate a ML model and Gibbs sampling and reconstruct the distribution of gas particles with temperature and 3D velocity. Our ML model is based on U-Net (Ronneberger et al., 2015), which is a supervised ML model originally developed for computer vision and consists of convolutional neural networks. The model learns gas dynamics of SN feedback in 3D from simulations. As described in Section 2.7, the ML model was trained with 300 simulations of an isolated SN in a molecular cloud with a total mass of $10^{6}$ $\mathrm{M_{\odot}}$ with a mass resolution of 1 $\mathrm{M_{\odot}}$ (Hirashima et al., 2023b). The ML model takes the 3D distribution ($64^{3}$ voxels with one side of 60 pc and a spatial resolution of $\sim 1$ pc) of density, temperature, and 3D velocities as an input and predicts those fields after the SN explosion by the time window $\tau=0.1$ Myr as an output, completed in Section 2.6. Since the time window is short enough compared to the local free fall time, we assume the cold gas particles in the voxels barely move, and the number of particles (the total mass) inside is preserved over the time window $\tau$. The model $\mathcal{M}$ is trained to learn the relation between input $X$ and output $y$, where $X$ and $y$ represent the distribution of physical quantities before and after the SN explosion in simulations, respectively, over the time window of 0.1 Myr. Suppose the trainable parameter $\theta$ and the predicted distribution $\hat{y}$, we summarize the procedure of the surrogate model as the following:
@@
Further details of our surrogate modeling are provided in Appendix B.
Figure 2 compares a direct simulation (the upper panels) by ASURA-FDPS and a simulation incorporating our surrogate model for SN feedback (the lower panel) by ASURA-FDPS-ML. The surrogate model reconstructs the asymmetric shell and hot region. The shell can evolve smoothly and stably even after the predicted particles return to the “Main” calculation domain. We also show the conversation of thermal energy and outer momentum in the reconstruction using our surrogate model, comparing it to a low-resolution simulation in Appendix B.4.
@@ 4 Results for an isolated dwarf galaxy simulation
Run Name Supernova Note
Thermal FB Surrogate Model
SN-DT Fiducial run
SN-ML
@@ 4.1 Morphology Comparison
Figure 3 shows the face-on surface density of the gas for two models summarized in Table 1, the fiducial run (SN-DT) on the left and the run with ML (SN-ML) on the right. We find little difference between the two runs when it comes to the morphological structure of the galactic disc. While the morphological structures that emerge on both simulations are supper bubbles driven by clustered feedback, we note that these volume-filling structures are not identical in the two reference runs we carried out. The reason for this is twofold. First of all, the binary executables of the code are not identical, which will lead to small-scale differences in execution. Second, some of the physics modules, such as the IMF sampling for star formation, introduce some randomness in the exact stellar population that is obtained, which in turn influences the feedback cycle due to the different lifetimes of massive stars of a given mass.
The ISM structures in our simulations are visually similar to other recent simulations of similar isolated dwarf galaxies (e.g., Smith et al., 2021; Deng et al., 2024; Gutcke et al., 2021; Hu, 2019). Fine mass and time resolutions are generally required for accurately resolving such superbubbles; otherwise, the hot gas within the bubble cools too quickly, a problem typically referred to as over-cooling. In our new scheme ASURA-FDPS-ML, however, despite the longer timesteps, several distinct superbubbles emerged.
Figure 3: Face-on surface density of gas at $t=100$ Myr for all models. Color represents column density in $\mathrm{10^{10}~{}M_{\odot}~{}kpc^{-2}}$. Left: SN-DT. Right: SN-ML. The different model variations shown in this Figure are summarized in Table 1.
@@ 4.2 Star Formation History
Figure 4 shows the star formation history (SFH), representing the SFRs averaged with an interval of 40 Myr as a function of time. For reference, the red dashed line shows the SFR obtained from the MFM simulation in Steinwandel & Goldberg (2023) as in their WLM-fid simulation. Since the IC of a disk dwarf galaxy does not have steller feedback initially, the galaxies have a high SFR from 0 to 80 Myr.
The simulations, SN-DT and SN-ML, have a similar SFR trend, they marginally differ; SN-DT shows an excess of star formation activity in the early stages of evolution, while SN-ML shows an excess of star formation activity in the late stages of the evolution. However, the agreement in the integrated star formation history (total stellar mass formed) is within a few percent between the two simulations, which is within so-called run-to-run variations that have been reported in previous studies of similar simulations (e.g., Steinwandel et al., 2023).
Figure 4: Star formation histories averaged with an interval of 40 Myr as a function of time for each run as listed in Table 1. The red dashed line shows the fiducial run in Steinwandel & Goldberg (2023).
@@ 4.4 Definition of Outflow Rates and Loading Factors
We follow the definition of the discrete outflow rates for mass and energy proposed in literature (Hu et al., 2017; Hu, 2019; Steinwandel & Goldberg, 2023) with the net total flow rates defined as $\dot{M}=\dot{M}_{\mathrm{out}}-\dot{M}_{\mathrm{in}}$ and $\dot{E}=\dot{E}_{\mathrm{out}}-\dot{E}_{\mathrm{in}}$. Outflows are computed as the sum of all gas with a positive or negative vertical velocity ($v_{z}$) in the slab located above or below the disc, respectively. Mass and energy outflow rates are defined as:
@@
We note that we find generally very good agreement with other studies in this field (Hu, 2019; Gutcke et al., 2021; Hislop et al., 2022; Steinwandel et al., 2023; Steinwandel & Goldberg, 2023; Steinwandel et al., 2024a; Porter et al., 2024) and we will discuss this in greater detail in Section 5.
@@ 4.5 Time Evolution and Phase Structure of the Outflows
Figure 6 shows the mass (top) and energy (bottom) outflow rates (left) and loading factors (right). The yellow lines represent the model SN-ML, and the blue lines represent the model SN-DT. The red dashed lines are the mean values obtained in the MFM simulation (Steinwandel & Goldberg, 2023) as in their WLM-fid simulation. We measure these flow rates and loading factors at two different heights, at 1 kpc (solid) and 10 kpc (dashed).
@@ 4.6 Transition of Sub-sonic-Supersonic in the Outflows
Figure 7: Two-dimensional joint PDFs of soundspeed and outflow velocity weighted by the mass outflow rate (flux) at height $z=1$ kpc (top) and at height $z=10$ kpc (bottom) for the models SN-DT (left) and SN-ML (right). The mass outflow rates are averaged between 200 Myr with an interval of 2 Myr. The gray dotted lines in the pdfs define contours equal to Bernoulli velocity defined by Equation (16)
@@ 4.7 Gas Density Distribution for SN Explosions
Figure 9 shows the probability distribution function of SNe as a function of the ambient density for the model SN-DT (blue) and the model SN-ML (yellow). Each physical model shows a bimodal distribution that peaks at a number density of $10^{-3}$ and $10^{3}$ $\mathrm{cm^{-3}}$, which is consistent with previous results of the SN-environmental density distribution of Hu et al. (2017), Gutcke et al. (2021) and Smith et al. (2021). The first peak around $10^{-3}~{}\mathrm{cm^{-3}}$ corresponds to the densities of typical super bubble interiors in those studies. The second peak around $10^{3}~{}\mathrm{cm^{-3}}$ is larger than the SF threshold and is indicative of the SN-environmental distribution without early stellar feedback as outlined by Hu et al. (2017) and Smith et al. (2021). We note that this peak at high SN-environmental densities is shifted to lower environmental densities by half a dex in the model SN-ML compared to the model SN-DT. This fact is indicative that SN-feedback remains ineffective in those dense environments in the resolved simulation SN-DT, and the ML algorithm in the simulation SN-ML can, in fact, circumvent this behavior.
@@ 5.1 Clustering of SNe
It has recently been pointed out that the clustering of SN-explosions plays a major role in the potential of a (dwarf) galaxy to drive an outflow. While there is still an ongoing debate on whether clustering (e.g., Emerick et al., 2018; Gutcke et al., 2021; Smith et al., 2021; Hislop et al., 2022; Hu et al., 2023; Steinwandel et al., 2023) vs. SN-environmental density (e.g., Hu et al., 2016, 2017; Steinwandel et al., 2020, 2024a, 2024b) is the primary driver of galactic outflows, there seems to be some consensus that they are “two sides of the same coin”. For instance, Hu et al. (2016, 2017), as well as the semi-analytical model, outlined that the lower environmental densities of SN-explosions are needed to generate a hot volume occupation fraction successfully launching the outflow. This intuition at the time has been built from the simulations of Hu et al. (2017), who, for the first time, invoked early stellar feedback which successfully quenched star formation on the cloud scale leading to lower star formation rates and thus lower outflow rates (but also loading factors) in comparison to simulations where SNe are the only feedback mechanism. While it is true that, generally speaking, the early stellar feedback lowers the environmental density, it also de-clusters star formation (e.g., Smith et al., 2021; Hu et al., 2023), which is considered to be the dominant factor. Theoretical models such as the one outlined in Naab & Ostriker (2017) and Steinwandel et al. (2020) make the argument that in order to establish a dominant volume filling phase one needs to have repeated explosion in gas with $n_{\mathrm{H}}<$ 0.1 cm-3, which is coincidentally the environmental density range that is reported with early stellar feedback in simulations such as Hu et al. (2017) and Smith et al. (2021). However, they ignore that these conditions can be generated in a strongly clustered (SN-only) perspective, which produces clearly stronger outflows based on numerical simulations.
@@
One possibility is that the simulation SN-DT, with its pure thermal feedback, may be subject to overcooling in such dense regions, suppressing the expansion of SN shells immediately after the explosion. Since the simulation SN-ML is bypassing this step with ML, one would never encounter such behavior leading to one prominent peak in explosions around 10-3 cm-3. There is some evidence for this since the total amount of SNe is around 13 percent for the model SN-DT and about 20 percent for the model SN-ML. Hence, subsequent SNe in the model SN-ML can (somewhat by construction) occur in lower-density environments. To what degree the ML framework helps to model more accurate thermal energy and momentum input in high-density regions in the simulation SN-ML remains somewhat hard to say as we find an undershoot of momentum in low-resolution simulations of isolated bubbles (see Appendix. B.4). It seems that the diffuse bubble can expand in high-density star-forming regions more efficiently, and as a result, short-lived stars may tend to explode in regions with much smaller densities than the SF threshold.
It is also interesting to compare SN-DT and SN-ML with similar models of a star formation efficiency of 2 percent without photo-ionization: noPE-noPI-SN in Hu et al. (2017), FIXED_SN_ENERGY in Gutcke et al. (2021), and SFE02noPI in Hislop et al. (2022). Their initial condition of the dwarf galaxy is similar settings as Steinwandel & Goldberg (2023). As for the implementation of SN feedback in Hislop et al. (2022), the thermal energy of SNe was injected at the ambient number density of $n_{\mathrm{H}}<10\>\mathrm{cm^{-3}}$ where individual SNe are well resolved and the expected outer momentum is input in dense regions (Hu et al., 2016; Steinwandel et al., 2020). The ambient density distributions of SN explosions for SFE02noPI also have a bimodal distribution with peaks around $10^{-4}$ and $10^{0}~{}\mathrm{cm^{-3}}$.
SN-ML gains more SNe around $10^{0}~{}\mathrm{cm^{-3}}$ compared to SN-DT. This can be because our ML model projects the prediction of the high resolution (1 $\mathrm{M_{\odot}}$) to a simulation with low resolution (4 $\mathrm{M_{\odot}}$). The large population of SNe explodes in the evacuated regions of previous SNe, which makes the feedback more efficient. After the first SN explodes on a gas cloud in our resolved simulations with a mass resolution of 1 $\mathrm{M_{\odot}}$, a diffused bubble can reach a density of $10^{-2}-10^{-4}$ $\mathrm{cm^{-3}}$, close values reported in Gutcke et al. (2021) and Hislop et al. (2022).

### v2

@@ 2.6 Hybrid on-the-fly ML-driven simulations
Figure 1 shows the schematic diagram of our presenting framework, ASURA-FDPS-ML. It has two kinds of MPI communicators: “Main” and “Pool” hereafter. A whole galaxy, including $N$-body and SPH, is directly computed in main processes using hundreds of MPI processes (“Main”), while SN feedback in dense regions ($>1~{}\mathrm{cm^{-3}}$) is handled by a surrogate model in “Pool” processes. As SNe in the less dense ambient gas are more kinetically resolved (Kim & Ostriker, 2015; Steinwandel et al., 2020; Gutcke et al., 2022a), when SNe explode in the environments at a hydrogen number density of $<1\mathrm{cm^{-3}}$, the energy of SNe ($10^{51}$ erg) is injected into the neighboring 100 SPH particles weighted using the Wendland $C^{4}$ kernel (Dehnen & Aly, 2012). Every timestep, “Main” checks if any SN occurs in dense regions ($>1~{}\mathrm{cm^{-3}}$) or not. If SNe occur, the ambient gas particles are sent to a “Pool” process. The surrogate model in the “Pool” process reconstructs the distribution of gas particles after a time window $\tau$ from the SN explosion. Once the “Main” processes are advanced for $\tau$, the predicted gas particles are returned to “Main” from “Pool”. If multiple SNe occur in a dense, compact region, our surrogate model handles the first explosion, which reduces the local gas density. Subsequent SNe then occur in a lower-density environment ($<1~{}\mathrm{cm^{-3}}$) and are automatically treated by the direct calculation method.
@@ 3 Test for our surrogate model of SN feedback
Figure 2: Comparison between simulations and ML for an isolated SN. Upper: The numerical simulation results. Lower: The numerical simulation results with our surrogate SN feedback model. Cross section of snapshots at $t=0,~{}10^{5},~{}\mathrm{and}~{}4\times 10^{5}$ years are listed up from left to right.
Our surrogate model is designed to incorporate a ML model and Gibbs sampling and reconstruct the distribution of gas particles with temperature and 3D velocity. Our ML model is based on U-Net (Ronneberger et al., 2015), which is a supervised ML model originally developed for computer vision and consists of convolutional neural networks. The model learns gas dynamics of SN feedback in 3D from simulations. As described in Section 2.7, the ML model was trained with 300 simulations of an isolated SN in a molecular cloud with a total mass of $10^{6}$ $\mathrm{M_{\odot}}$ with a mass resolution of 1 $\mathrm{M_{\odot}}$ (Hirashima et al., 2023b). The ML model takes the 3D distribution ($64^{3}$ voxels with one side of 60 pc and a spatial resolution of $\sim 1$ pc) of density, temperature, and 3D velocities as an input and predicts those fields after the SN explosion by the time window $\tau=0.1$ Myr as an output, completed in Section 2.6. Since the time window is short enough compared to the local free-fall time, we assume the cold gas particles in the voxels barely move, and the number of particles (the total mass) inside is preserved over the time window $\tau$. The model $\mathcal{M}$ is trained to learn the relation between input $X$ and output $y$, where $X$ and $y$ represent the distribution of physical quantities before and after the SN explosion in simulations, respectively, over the time window of 0.1 Myr. Suppose the trainable parameter $\theta$ and the predicted distribution $\hat{y}$, we summarize the procedure of the surrogate model as the following:
@@
Further details of our surrogate modeling are provided in Appendix C.
Figure 2 compares a direct simulation (the upper panels) by ASURA-FDPS and a simulation incorporating our surrogate model for SN feedback (the lower panel) by ASURA-FDPS-ML. The surrogate model reconstructs the asymmetric shell and hot region. The shell can evolve smoothly and stably even after the predicted particles return to the “Main” calculation domain. We also show the conversation of thermal energy and outer momentum in the reconstruction using our surrogate model, comparing it to a low-resolution simulation in Appendix C.4.
@@ 4 Results for an isolated dwarf galaxy simulation
Run Name Supernova Feedback Time-Step Note
SN-DT Thermal Feedback Determined by eq. (1) Fiducial run
SN-ML Thermal Feedback (if $n_{\mathrm{H}}<1\mathrm{cm^{-3}}$) Fixed 2,000 yr
Our Surrote Modeling (if $n_{\mathrm{H}}>1\mathrm{cm^{-3}}$)
@@ 4.1 Morphology Comparison
Figure 3 shows the face-on surface density of the gas for two models summarized in Table 1, the fiducial run (SN-DT) on the left and the run with ML (SN-ML) on the right. We find little difference between the two runs when it comes to the morphological structure of the galactic disc. While the morphological structures that emerge on both simulations are super bubbles driven by clustered feedback, we note that these volume-filling structures are not identical in the two reference runs we carried out. The reason for this is twofold. First of all, the binary executables of the code are not identical, which will lead to small-scale differences in execution. Second, some of the physics modules, such as the IMF sampling for star formation, introduce some randomness in the exact stellar population that is obtained, which in turn influences the feedback cycle due to the different lifetimes of massive stars of a given mass.
The ISM structures in our simulations are visually similar to other recent simulations of similar isolated dwarf galaxies (e.g., Smith et al., 2021; Deng et al., 2024; Gutcke et al., 2022a; Hu, 2019). Fine mass and time resolutions are generally required for accurately resolving such superbubbles; otherwise, the hot gas within the bubble cools too quickly, a problem typically referred to as numerical over-cooling (e.g., Pearce et al., 1999; Thacker et al., 2000; Croft et al., 2001; Springel & Hernquist, 2002). In our new scheme ASURA-FDPS-ML, however, despite the longer timesteps, several distinct superbubbles emerged.
Figure 3: Face-on surface density of gas at $t=100$ Myr for all models. Color represents column density in $\mathrm{10^{10}~{}M_{\odot}~{}kpc^{-2}}$. Left: SN-DT. Right: SN-ML. The different model variations shown in this Figure are summarized in Table 1. An animation showing the time evolution of the face-on gas surface density is available. It spans 300 Myr of simulation time and lasts 20 seconds in real time.
@@ 4.2 Star Formation History
Figure 4 shows the star formation history (SFH), representing the SFRs averaged with an interval of 10 Myr as a function of time. For reference, the red dashed line shows the SFR obtained from the MFM simulation in Steinwandel & Goldberg (2025) as in their WLM-fid simulation. Since the IC of a disk dwarf galaxy does not have steller feedback initially, the galaxies have a high SFR from 0 to 80 Myr. The simulations, SN-DT and SN-ML, have a similar SFR trend, they marginally differ; SN-DT shows an excess of star formation activity in the early stages of evolution, while SN-ML shows an excess of star formation activity in the late stages of the evolution. However, the agreement in the integrated star formation history (total stellar mass formed) is within a few percent between the two simulations, which is within so-called run-to-run variations that have been reported in previous studies of similar simulations (e.g., Steinwandel et al., 2023).
Figure 4: Star formation histories averaged with an interval of 10 Myr as a function of time for each run as listed in Table 1. The red dashed line shows the fiducial run in Steinwandel & Goldberg (2025).
@@ 4.4 Definition of Outflow Rates and Loading Factors
We follow the definition of the discrete outflow rates for mass and energy proposed in literature (Hu et al., 2017; Hu, 2019; Steinwandel & Goldberg, 2025) with the net total flow rates defined as $\dot{M}=\dot{M}_{\mathrm{out}}-\dot{M}_{\mathrm{in}}$ and $\dot{E}=\dot{E}_{\mathrm{out}}-\dot{E}_{\mathrm{in}}$. Outflows are computed as the sum of all gas with a positive or negative vertical velocity ($v_{z}$) in the slab located above or below the disc, respectively. Mass and energy outflow rates are defined as:
@@
We note that we find generally very good agreement with other studies in this field (Hu, 2019; Gutcke et al., 2022a; Hislop et al., 2022; Steinwandel et al., 2023; Steinwandel & Goldberg, 2025; Steinwandel et al., 2024a; Porter et al., 2024) and we will discuss this in greater detail in Section 5.
@@ 4.5 Time Evolution and Phase Structure of the Outflows
Figure 6 shows the mass (top) and energy (bottom) outflow rates (left) and loading factors (right). The yellow lines represent the model SN-ML, and the blue lines represent the model SN-DT. The red dashed lines are the mean values obtained in the MFM simulation (Steinwandel & Goldberg, 2025) as in their WLM-fid simulation. We measure these flow rates and loading factors at two different heights, at 1 kpc (solid) and 10 kpc (dashed).
@@ 4.6 Transition of Sub-sonic-Supersonic in the Outflows
Figure 7: Two-dimensional joint PDFs of soundspeed and outflow velocity weighted by the mass outflow rate (flux) at height $z=1$ kpc (top) and at height $z=10$ kpc (bottom) for the models SN-DT (left) and SN-ML (right). The mass outflow rates are averaged between 200 Myr with an interval of 2 Myr. The gray dotted lines in the pdfs define contours equal to Bernoulli velocity defined by Equation (16).
@@ 4.7 Gas Density Distribution for SN Explosions
Figure 9 shows the probability distribution function of SNe as a function of the ambient density for the model SN-DT (blue) and the model SN-ML (yellow). Each physical model shows a bimodal distribution that peaks at a number density of $10^{-3}$ and $10^{3}$ $\mathrm{cm^{-3}}$, which is consistent with previous results of the SN-environmental density distribution of Hu et al. (2017), Gutcke et al. (2022a) and Smith et al. (2021). The first peak around $10^{-3}~{}\mathrm{cm^{-3}}$ corresponds to the densities of typical super bubble interiors in those studies. The second peak around $10^{3}~{}\mathrm{cm^{-3}}$ is larger than the SF threshold and is indicative of the SN-environmental distribution without early stellar feedback as outlined by Hu et al. (2017) and Smith et al. (2021). We note that this peak at high SN-environmental densities is shifted to lower environmental densities by half a dex in the model SN-ML compared to the model SN-DT. This fact is indicative that SN-feedback remains ineffective in those dense environments in the resolved simulation SN-DT, and the ML algorithm in the simulation SN-ML can, in fact, circumvent this behavior.
@@ 5.1 Clustering of SNe
It has recently been pointed out that the clustering of SN-explosions plays a major role in the potential of a (dwarf) galaxy to drive an outflow. While there is still an ongoing debate on whether clustering (e.g., Emerick et al., 2018; Gutcke et al., 2022a; Smith et al., 2021; Hislop et al., 2022; Hu et al., 2023; Steinwandel et al., 2023) vs. SN-environmental density (e.g., Hu et al., 2016, 2017; Steinwandel et al., 2020, 2024a, 2024b) is the primary driver of galactic outflows, there seems to be some consensus that they are “two sides of the same coin”. For instance, Hu et al. (2016, 2017), as well as the semi-analytical model, outlined that the lower environmental densities of SN-explosions are needed to generate a hot volume occupation fraction successfully launching the outflow. This intuition at the time has been built from the simulations of Hu et al. (2017), who, for the first time, invoked early stellar feedback which successfully quenched star formation on the cloud scale leading to lower star formation rates and thus lower outflow rates (but also loading factors) in comparison to simulations where SNe are the only feedback mechanism. While it is true that, generally speaking, the early stellar feedback lowers the environmental density, it also de-clusters star formation (e.g., Smith et al., 2021; Hu et al., 2023), which is considered to be the dominant factor. Theoretical models such as the one outlined in Naab & Ostriker (2017) and Steinwandel et al. (2020) make the argument that in order to establish a dominant volume filling phase one needs to have repeated explosion in gas with $n_{\mathrm{H}}<$ 0.1 cm-3, which is coincidentally the environmental density range that is reported with early stellar feedback in simulations such as Hu et al. (2017) and Smith et al. (2021). However, they ignore that these conditions can be generated in a strongly clustered (SN-only) perspective, which produces clearly stronger outflows based on numerical simulations.
@@
One possibility is that the simulation SN-DT, with its pure thermal feedback, may be subject to overcooling in such dense regions, suppressing the expansion of SN shells immediately after the explosion. Since the simulation SN-ML is bypassing this step with ML, one would never encounter such behavior leading to one prominent peak in explosions around 10-3 cm-3. There is some evidence for this since the total amount of SNe is around 13 percent for the model SN-DT and about 20 percent for the model SN-ML. Hence, subsequent SNe in the model SN-ML can (somewhat by construction) occur in lower-density environments. To what degree the ML framework helps to model more accurate thermal energy and momentum input in high-density regions in the simulation SN-ML remains somewhat hard to say as we find an undershoot of momentum in low-resolution simulations of isolated bubbles (see Appendix. C.4). It seems that the diffuse bubble can expand in high-density star-forming regions more efficiently, and as a result, short-lived stars may tend to explode in regions with much smaller densities than the SF threshold.
It is also interesting to compare SN-DT and SN-ML with similar models of a star formation efficiency of 2 percent per free-fall time without photo-ionization: noPE-noPI-SN in Hu et al. (2017), FIXED_SN_ENERGY in Gutcke et al. (2022a), and SFE02noPI in Hislop et al. (2022). Their initial condition of the dwarf galaxy is similar settings as Steinwandel & Goldberg (2025). In the implementation of SN feedback in Hislop et al. (2022), the thermal energy of SNe was injected at the ambient number density of $n_{\mathrm{H}}<10\>\mathrm{cm^{-3}}$ where individual SNe are well resolved and the expected outer momentum is input in dense regions (Hu et al., 2016; Steinwandel et al., 2020). The ambient density distributions of SN explosions for SFE02noPI also have a bimodal distribution with peaks around $10^{-4}$ and $10^{0}~{}\mathrm{cm^{-3}}$.
SN-ML gains more SNe around $10^{0}~{}\mathrm{cm^{-3}}$ compared to SN-DT. This can be because our ML model projects the prediction of the high resolution (1 $\mathrm{M_{\odot}}$) to a simulation with low resolution (4 $\mathrm{M_{\odot}}$). The large population of SNe explodes in the evacuated regions of previous SNe, which makes the feedback more efficient. After the first SN explodes on a gas cloud in our resolved simulations with a mass resolution of 1 $\mathrm{M_{\odot}}$, a diffused bubble can reach a density of $10^{-2}-10^{-4}$ $\mathrm{cm^{-3}}$, close values reported in Gutcke et al. (2022a) and Hislop et al. (2022).